OpenStudy (anonymous):

Find LCD for each list of rational expressions. 2x/3x^2 + 4X +1 , 7/2x^2 - x - 1 I solved 2x/3x^2 + 4X +1 all fine. 3x(x+1) + 1(x+1) (3x+1)(x+1) I get stuck on 7/2x^2 - x - 1

OpenStudy (neer2890):

how can you say that \[\frac{ 2x }{ 3x ^{2} }+4x+1=(3x+1)(x+1)\]

OpenStudy (anonymous):

we are ignoring the numerator in this problem

OpenStudy (neer2890):

I think it's \[\frac{ 2x }{ 3x ^{2}+4x+1 }\]

OpenStudy (anonymous):

3x2+4x+1 = 3x^2 + 3x +x +1 3x(x+1) +1(x+1) (3x+1)(x-1)

OpenStudy (neer2890):

\[2x ^{2}-x-1\] \[=2x ^{2}-2x+x-1\] =2x(x-1)+1(x-1) =(2x+1)(x-1)

OpenStudy (anonymous):

see this is where i am confused

OpenStudy (anonymous):

sorry! i ment 7/2x^2 - 2x - 1

OpenStudy (anonymous):

for the original equation

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

your right

OpenStudy (anonymous):

i was reading problem wrong

OpenStudy (neer2890):

\[\frac{ 7 }{ 2x ^{2}-x+1 }=\frac{ 7 }{ (2x+1)(x-1) }\]

OpenStudy (neer2890):

I think you got it...!